twice the sum of a number and 5 is equal to three times the diffrence of the number and 7
Answer by ewatrrr(24785)

twice the sum of a number and 5 is equal to three times the diffrence of the number and 7
Write as You Read ( let n represent the "number)
2(n+5) = 3(n-7)
Solve
2n+10 = 3n - 21
31 = n
AND...
2*36 = 3(24) = 72 CHECKS!